OpenStudy (anonymous):

Please help! Rewrite the expression as a single logarithm log A. The the function A= 4log(x)-3log(x^+1)+3log(x-1)

OpenStudy (anonymous):

please see the question what does ^ mean ?

OpenStudy (anonymous):

I don't understand it

OpenStudy (anonymous):

Sorry I meant to put ^2

OpenStudy (anonymous):

\[\log x^4-\log \left( x^2+1 \right)^3+\log \left( x-1 \right)^3\] \[\log a+logb-\log c-\log d+loge=\log \frac{ abe }{ cd }\]